If seven matrons assembled for a meeting shake hands with one another,how many handshakes take place
Answer by stanbon(75887)

If seven matrons assembled for a meeting shake hands with one another,how many handshakes take place
-------
Ans: 6+5+4+3+2+1 = 3*7 = 21
-------------------------------
Cheers,
Stan H.